OpenStudy (anonymous):

Forces with magnitudes of 70 pounds, 40 pounds, and 60 pounds act on an object at angles of -30, 45, and 135, respectively, with the positive x axis. Find the direction of the magnitude of the resultant of these forces

OpenStudy (anonymous):

The resultant x-component is: 70cos(- 60) + 40cos(45) + 60cos(120) = 33.284 pounds. The resultant y-component is: 70sin(- 60) + 40sin(45) + 60sin(120) = 19.624 pounds. The resultant is: sqrt(33.284^2 + 19.624^2) = 38.6 pounds to 3 sig. fig. If t is its direction measured from the positive x-axis: tan(t) = 19.624 / 33.284 Both x- and y-components are positive, indicating a quadrant 1 angle: t = 30.5 deg.

OpenStudy (anonymous):

This has incorrect points

OpenStudy (anonymous):

isn't this physics?

OpenStudy (anonymous):

then again, physics is maths

OpenStudy (anonymous):

I'm looking for -30, 45, and 135

OpenStudy (anonymous):

there's too little data for the previous question ket-kat

OpenStudy (anonymous):

There's enough, I just need to know how to set it up